OpenStudy (anonymous):

A woman is selling peanuts and cashews, cashews for 8 dollars a pound and peanuts for 3 dollars a pound. She wants to make a mixture that costs 4 dollars a pound. If she uses 10 pounds of peanuts, how much will the total mixture weigh?

OpenStudy (anonymous):

solve \[30+8x=4(10+x)\] for x

OpenStudy (amistre64):

why is first place "pee yellow"?

OpenStudy (anonymous):

reasoning that the 10 pounds of peanuts are $3 a pound and the cashews are $8 and the total will be 10 + x

OpenStudy (anonymous):

color changes constantly

OpenStudy (anonymous):

sometimes yellow, sometimes grayed out, and different shades of red ishness

OpenStudy (anonymous):

i miss the jiggle effect though. what happened to that???

OpenStudy (anonymous):

mixture will weight 12.5 pounds

OpenStudy (anonymous):

I know the answer is 12.5, but can someone show me the right equation? If the top one is used, x=-2.5 and the total mixture wouldn't weigh 12.5, right?

OpenStudy (anonymous):

lets suppose in 1 pound mixture x kg is peanut and 1-x kg is cashew so x*3+(1-x)*8=4 by solving we get x=4/5 that means 4/5 kg of peanuts and 1/5 kg of cashew that means 4 times more peanuts then cashew so while there is 10 kg of peanut then there will be 10/4 kg of cashew so 10+10/4=12.5pounds